solve.
sqrt(x-1)=x-3

Answer by stanbon(75887)

solve.
sqrt(x-1)=x-3
---------------
Square both sides to get:
x-1 = x^2-6x+9
x^2-7x+10=0
(x-5)(x-2)=0
x = 5 or x=2
-----------------
Check the answers in the original equation:
Does x=5?
sqrt(4)=5-3
True
--------------
Does x=2?
sqrt(1) = 2-3
sqrt1 = -1
False
--------------
Only solution: x=5
======================
Cheers,
Stan H.